# A Cascaded Random Access Quantum Memory

### v1

@@ Appendix B System Hamiltonian
Figure 6: (a) RAQM circuits diagram. For simplicity, we only show the coupling to the buffer mode $B_{i}$ and the storage mode $S_{j}$. The coupler is capacitively coupled to all buffer, storage, and dump modes. (b) Experimentally measured $B_{1}S_{j}$ sideband rates.
@@
Fig. 6(a) shows the RAQM circuit diagram. The effective capacitance and inductance for different modes are calculated from Black-box quantization [38] using the admittance extracted through HFSS simulations. To calculate the memory sideband rate analytically, we first consider the subsystem containing $B_{i}$, coupler, and $S_{j}$ for simplicity. The corresponding dressed frequencies are $\omega_{bi}$, $\omega_{c}$, and $\omega_{sj}$. We assume that junctions on the coupler SQUID are symmetric with Josephson energy $E_{Jc}$. The coupler SQUID potential $U$ is:
$U=E_{Jc}\cos\left(\varphi_{c}+\frac{\pi\Phi_{\text{ext}}}{\Phi_{0}}\right).$ (2)
@@
Here,$\Phi_{\text{ext}}$ is the external flux threading the SQUID loop and $\varphi_{c}$ is the phase variable associated with the coupler mode in the dressed basis. Under the external RF flux modulation of amplitude $\epsilon$ and at frequency $\omega_{d}$ as well as a continuous DC flux bias $\varphi_{\text{DC}}$:
$\frac{\pi\Phi_{\text{ext}}}{\Phi_{0}}=\varphi_{\text{DC}}+\varepsilon\sin(\omega_{d}t),$ (3)
We can rewrite the coupler potential as
$U=E_{Jc}\cos\left(\varphi_{c}+\varphi_{\text{DC}}+\varepsilon\sin(\omega_{d}t)\right).$ (4)
@@
Assuming $\varepsilon$ to be small ($\varepsilon=0.035\ll 1$ from simulations) , we can expand Eq. 4 up to first order in $\varepsilon$ as:
$U\approx E_{Jc}(\cos(\varphi_{c}+\varphi_{\text{DC}})+\varepsilon\sin(\varphi_{c}+\varphi_{\text{DC}})\sin(\omega_{d}t))$ (5)
@@
When the coupler is dispersively coupled to the other modes, in the appropriate rotating frame, $\varphi_{c}$ can be approximated using annihilation operators:
$\varphi_{c}$ $=\varphi_{\text{zpf}}(a_{c}e^{i\omega_{c}t}-\frac{g_{cbi}}{\omega_{c}-\omega_{bi}}a_{bi}e^{i\omega_{bi}t}$
$-\frac{g_{csj}}{\omega_{c}-\omega_{sj}}a_{sj}e^{i\omega_{sj}t}+h.c.)$ (6)
@@
where $\varphi_{\text{zpf}}$ is the zero point fluctuation of the phase variable. Here the dressed operators for the coupler mode include contributions from each mode, normalized by participation factors of the form $g/\Delta$ where $g$ is the coupling strength and $\Delta$ is the detuning between the coupler and a given mode. We can plug Eq. 6 into Eq. 5 and expand the potential up to the second order in the participation factors. We choose $\omega_{d}=\omega_{bi}-\omega_{sj}$ to activate the beamsplitter interactions. Keeping only the time-independent terms, we have:
@@
$U\approx\varepsilon E_{Jc}\sin(\varphi_{\text{DC}})\frac{g_{cbi}g_{csj}}{2(\omega_{c}-\omega_{bi})(\omega_{c}-\omega_{sj})}(a_{bi}a_{sj}^{\dagger}+h.c.),$ (7)
@@
which describes the beamsplitter rate. In experiments, we calibrate this beamsplitter gate by preparing a photon in the buffer mode and activating the beamsplitter drive for $\sim 100\mathrm{\mu}$s (equivalent to $\sim 50$ swaps) for a range of frequencies. We choose the resonant frequency which maximizes the probability of finding the photon in the storage mode. The repeated sequence of swap gates helps us calibrate the pulse parameters in presence of coherent errors. For calibration in presence of bichromatic drive tones or strong stark shifts, we refer the reader to ref. [27]. Fig. 6(b) shows the sideband rates used in the RAQM experiments.
@@
The cross-Kerr $\chi_{b_{i}s_{j}}$ between buffer mode $B_{i}$ and storage mode $S_{j}$, and the modes’ self-Kerrs $\{k_{bi},k_{sj}\}$ can be calculated using the effective mode capacitance $\{C_{bi},C_{sj}\}$ and inductance $\{L_{bi},L_{sj}\}$ [38]:
$\chi_{b_{i}s_{j}}$ $=-2\sqrt{k_{bi}k_{sj}},$ (8)
$k_{bi(sj)}$ $=\frac{L_{bi(sj)}}{C_{bi(sj)}}\frac{C_{c}}{L_{c}}E_{C_{c}},$ (9)
@@ Appendix C Active Reset
The long coherence of the storage modes results in an extended system reset time ($55T_{1}>$ 5 ms). To speed up all the experiments, we implement an active reset protocol modularly, including buffer mode reset, qubit reset, and storage mode reset.
Figure 7 (a) illustrates the sideband interactions between the buffer and dump modes. We separately prepare $\left|1\right\rangle$ in $B_{1}$ and $B_{2}$, then activate the coupler modulation at the difference frequencies of $B_{1}D_{1}$ and $B_{2}D_{2}$ for a variable time. The lossy dump modes, acting as cold reservoirs, autonomously dissipate the received excitations, effectively cooling the buffer modes. The residual population in the buffer modes is swapped back to the qubit’s $\left|f\right\rangle$ level, rotated to $\left|e\right\rangle$ level, and subsequently measured. The residual population of both buffer modes remains constant at $<10\%$ after $12\,\mu\mathrm{s}$ of sideband interactions.
Figure 7 (b) illustrates our RAQM active reset protocol, performed before the main experiments. The buffer modes are evacuated first. In the memory experiments, both the buffer modes and the qubit theoretically use, at most, the first excited states. Consequently, the protocol ensures state reset upto the second excitation for all modes. Since the qubit-$B_{i}$ dispersive shift $\chi_{qb_{i}}$ is comparable to the sideband rate, we sequentially choose $9$ different coupler RF modulations frequencies for each $B_{i}D_{i}$ sideband, accounting for the dressing effects of the dispersive shifts. After evacuating both buffer modes, we sequentially reset the qubit’s $\left|e\right\rangle$ and $\left|f\right\rangle$ states. First, we measure the qubit state using a $1\,\mu\mathrm{s}$ readout pulse, wait for $2\,\mu\mathrm{s}$, and conditionally flip the qubit state if it is in the $\left|e\right\rangle$ state. Next, we apply a qubit $\left|f\right\rangle$-$\left|e\right\rangle$ $\pi$ pulse and repeat the active reset for the $\left|e\right\rangle$ state. To reset the storage modes, we sequentially apply $7$ swap pairs, transferring the population from $S_{i}$ to $B_{1}$, then dump the population from $B_{1}$ to $D_{1}$ and autonomously evacuated. Finally, before starting the memory experiments, we idle the system for $75\,\mu\mathrm{s}$ so that any remaining low-coherence modes are fully reset. The experiments are initiated only when the qubit is confirmed in the $\left|g\right\rangle$. The qubit readout performance in our experiments is shown in Table 4.
Figure 7: RAQM active reset. (a) Buffer-Dump modes swap. Dash line ($12\mu$s) shows the time experiments used to evacuate buffer modes. (b) Active reset sequence. The sequence starts with evacuating buffer modes. Multiple buffer-dump swaps at different frequencies (dressed by $\chi_{qb_{i}}$) are sequentially applied. Qubit state (up to $\left|f\right\rangle$ level) is actively reset. All Storage modes’ populations are sequentially swapped into $B_{1}$, then evacuated through $D_{1}$. After $75\,\mu\mathrm{s}$’ idling period, the experiment starts when the qubit state is $\left|g\right\rangle$.
@@ Appendix D System Temperature
We benchmark the cavity’s temperature using the Ramsey parity measurement [41]. To detect the thermal photon population in $B_{1}$, we repeatedly measure the parity information 80 times in a single experiment. Each measurement sequence consists of the following pulses: a transmon $\pi/2$ rotation, an idle period of $\pi/\chi$, another transmon $\pi/2$ rotation, a readout pulse ($1\,\mu\mathrm{s}$), and a readout evacuation period ($2.5\,\mu\mathrm{s}$). Each experiment produces an 80-bit string indicating the transmon’s state. Fig. 8(a) illustrates the string behavior when $B_{1}$ is initialized to $\left|1\right\rangle$.
Given a specific type of transmon state string, we quantify whether the cavity is populated by applying the hidden Markov process [41] to calculate the cavity population ratio:$\frac{P_{1}}{P_{0}}$. Here, the cavity state probabilities $P_{0}$ and $P_{1}$ correspond to the cavity being in $\left|0\right\rangle$ and $\left|1\right\rangle$, respectively. These probabilities are reconstructed using the backward propagation algorithm based on the transmon state string. Starting from the beginning of the state string, the algorithm iteratively calculates the likelihood of observing the measured readout signal up to the current position, assuming the cavity was initially in $\left|0\right\rangle$ or $\left|1\right\rangle$. Each transmon state string is then converted into a probability string.
@@
In Fig. 8(b), we analyzed $10^{5}$ measurements taken when the system is in its thermal equilibrium state. A higher ratio in the plot indicates a greater likelihood of photon population in the buffer modes. As the threshold increases (dashed line), measurement errors are filtered out, isolating events corresponding to thermal photons.
After extracting the thermal photon events, we need to compensate for two key factors to avoid overestimating the system temperature: detection efficiency and event assignment. The detection efficiency $\eta_{i}$ describes the probability of detecting a photon when $B_{i}$ is in $\left|1\right\rangle$. In our experiments, we separately prepare $\left|1\right\rangle$ in $B_{1}$ and $B_{2}$, perform the Ramsey parity measurement to count cavity photon events, and approximate $\eta_{i}$ using $2\times 10^{5}$ experiments. In Fig. 8(c), fewer photon events are captured as the threshold increases. Since both $B_{1}$ and $B_{2}$ are coupled to the transmon, thermal population in either cavity can trigger the detection. We assume that both modes have the same temperature $T_{c}$. Given the experimentally measured total photon counts $\delta$ and efficiency $\eta_{i}$, the temperature $T_{c}$ required to explain the counts should satisfy:
$\sum_{i=1,2}\frac{\eta_{i}}{-1+\exp{(\frac{\hbar\omega_{bi}}{k_{B}T_{c}}})}=\delta$ (11)
Fig. 8(d) shows the corrected thermal photon counts in $B_{1}$ and the corresponding temperature. We choose $10^{13}$ as our threshold ($\eta_{1}=46.6\%\pm 0.2\%$, $\eta_{2}=33.0\%\pm 0.2\%$) and list the system temperature (steady state temperature $T_{SS}$ and photon counts $\bar{n}_{ss}$, active reset temperature $T_{ar}$ and photon counts $\bar{n}_{ar}$ in Table 5. Qubit temperature is calculated by measuring $\left|e\right\rangle\leftrightarrow\left|f\right\rangle$ Rabi oscillation amplitude with and without initial $\pi$ pulse between $\left|g\right\rangle$ and $\left|e\right\rangle$. The temperature of the other storage and dump modes is measured with an additional $\pi$ swap between corresponding buffer modes initially. The coupler’s temperature is higher after the active reset, which can be mitigated by adding additional coupler reset through the dump modes or waiting longer ($>75\,\mu$s) before the experiments start.
@@ Appendix E Phase Calibration for RAQM experiments
Figure 9: RAQM Phase calibration. Each transmon gate’s virtual-$Z$ phase comprises three components: the active-access phase $\theta_{ii}$, the inactive-access phase $\theta_{ij}$ accumulated during storage read and write, and the time-dependent idle phase $\theta_{\Delta_{i}}(t)$ caused by hardware frequency tracking. (a) Pulse sequences for state transfer between transmon and $B_{1}$. (b) active-access phase $\theta_{ii}$ calibration for $S_{i}$. (c) Idle phase calibration extracts the Stark-shift $\Delta_{i}$ to $S_{i}$ through the Ramsey-like experiments. (d) Inactive-access phase $\theta_{ij}$ calibration on $S_{i}$ after $S_{j}$ access operations. $\theta_{ij}$ are calculated based on the maximum contrast phase $\tilde{\theta}_{i}j$ and $\Delta_{i}$. To improve calibration accuracy, $N+1$ writes and $N$ reads are repeated.
@@
Fig. 9 shows the process of virtual-$Z$ phase calibration. For simplicity, state read and write operations between the transmon and $B_{1}$ are represented using circuit symbols in Fig. 9(a). Fig. 9(b) shows the circuit to calibrate $\theta_{ii}$. The phase of the second transmon gate is swept, and $\theta_{ii}$ is determined as the phase that maximizes the final $\left|e\right\rangle$ population. Fig. 9(c) shows the circuit to calibrate $\theta_{ii}$. By sweeping the waiting time between two $B_{1}S_{i}$ swaps, we fit the Ramsey signal and extract the Stark-shift frequency $\Delta_{i}$. Fig. 9(d) shows the circuit to calibrate $\theta_{ij}$. Since the inactive-access phase is small, we repeat the read and write operations on $S_{j}$ for $5$ times. The maximum contrast phase $\tilde{\theta}_{ij}$, determined by the maximum final $\left|e\right\rangle$ population in the transmon, can be used to calculate $\theta_{ij}$:

### v2

@@ Appendix B System Hamiltonian
Figure 6: (a) False-color device image. (b) RAQM circuit diagram. For simplicity, only the couplings to the buffer mode $B_{i}$ and the storage mode $S_{j}$ are shown. The coupler is capacitively coupled to all buffer, storage, and dump modes. (c) Experimentally measured $B_{1}S_{j}$ sideband rates.
Supplementary Figure 6(a) provides a physical illustration of the RAQM. The dump modes are technically unprotected storage modes at higher frequencies that occupy the same locations as the storage modes. The other RAQM components are color-coded according to their physical placement.
@@
Supplementary Figure 6(b) shows the RAQM circuit diagram. The effective capacitances and inductances of the different modes are calculated using Black-box quantization [47] from the admittance extracted via HFSS simulations. To calculate the memory sideband rate analytically, we first consider the subsystem consisting of $B_{i}$, the coupler, and $S_{j}$ for simplicity. The corresponding dressed frequencies are $\omega_{b_{i}}$, $\omega_{c}$, and $\omega_{s_{j}}$. Assuming symmetric Josephson junctions with energy $E_{J}$, and a purely differential RF flux drive [39], the potential energy of the coupler SQUID is
$U=2E_{J}\cos\!\left(\varphi_{\text{ext}}\right)\cos\!\left(\varphi_{c}\right).$ (3)
@@
Here $\varphi_{\text{ext}}=\pi\Phi_{\text{ext}}/\Phi_{0}$ is the external flux threading the SQUID loop, and $\varphi_{c}$ is the phase variable associated with the coupler mode in the dressed basis. Under an external RF flux modulation of amplitude $\varepsilon$ and frequency $\omega_{d}$, together with a DC flux bias $\varphi_{\text{DC}}$, we have
$\varphi_{\text{ext}}=\varphi_{\text{DC}}+\varepsilon\sin(\omega_{d}t).$ (4)
@@
Assuming $\varepsilon\ll 1$ ($\varepsilon=0.035$ from simulations), we substitute Eq. 4 into Eq. 3 and expand to first order in $\varepsilon$:
$U\approx 2E_{J}\Bigl(\cos\varphi_{\text{DC}}-\varepsilon\sin(\omega_{d}t)\sin\varphi_{\text{DC}}\Bigr)\cos\varphi_{c}.$ (5)
@@
When the coupler is dispersively coupled to the other modes, in the appropriate rotating frame, $\varphi_{c}$ can be approximated by annihilation operators:
$\varphi_{c}$ $=\varphi_{\text{zpf}}\Bigl(a_{c}e^{i\omega_{c}t}-\frac{g_{cb_{i}}}{\Delta_{cb_{i}}}a_{b_{i}}e^{i\omega_{b_{i}}t}$
$\hskip 56.9055pt-\frac{g_{cs_{j}}}{\Delta_{cs_{j}}}a_{s_{j}}e^{i\omega_{s_{j}}t}+\text{h.c.}\Bigr),$ (6)
@@
where $\varphi_{\text{zpf}}$ is the zero-point fluctuation of the phase variable. The dressed operators for the coupler mode include contributions from each mode, normalized by participation factors of the form $g/\Delta$, where $g_{cx}$ is the coupling strength and $\Delta_{cx}=\omega_{c}-\omega_{x}$ is the detuning between the coupler and mode $x$.
Plugging Eq. 6 into Eq. 5 and expanding the potential up to second order in the participation factors, we choose $\omega_{d}=\omega_{b_{i}}-\omega_{s_{j}}$ to activate the beamsplitter interactions. Keeping only the time-independent terms, we obtain
@@
$U_{\text{swap}}\approx\underbrace{\varepsilon E_{J}\sin(\varphi_{\text{DC}})\varphi_{\text{zpf}}^{2}\frac{g_{cb_{i}}}{\Delta_{cb_{i}}}\frac{g_{cs_{j}}}{\Delta_{cs_{j}}}}_{g_{\textrm{swap},b_{i}s_{j}}}\bigl(a_{b_{i}}a_{s_{j}}^{\dagger}+\text{h.c.}\bigr),$ (7)
@@
where $g_{\textrm{swap},b_{i}s_{j}}$ denotes the beamsplitter rate. In experiments, we calibrate this beamsplitter gate by preparing a photon in the buffer mode and activating the beamsplitter drive for $\sim 100\,\mu\text{s}$ (equivalent to $\sim 50$ swaps) across a range of frequencies. We then choose the resonant frequency that maximizes the probability of finding the photon in the storage mode. The repeated sequence of swap gates helps calibrate the pulse parameters in the presence of coherent errors. For calibration under bichromatic drive tones or strong Stark shifts, we refer the reader to Ref. [39].
Supplementary Figure 6(c) shows the sideband rates used in the RAQM experiments. The primary limitation to our gate speeds are frequency collisions between these sidebands. For example, the $B_{1}S_{1}$ ($\omega_{s_{1}}-\omega_{b_{1}}=2\pi\times 349\,\text{MHz}$) and the $S_{1}S_{3}$ ($\omega_{s_{3}}-\omega_{s_{1}}=2\pi\times 348\,\text{MHz}$) differ by only $1\,\text{MHz}$. The size of collisions for other $B_{1}S_{j}$ sidebands ranges from $2{-}20\,\text{MHz}$. These frequency collisions limit our gate speeds to $300{-}500\,\text{kHz}$. Such issues can be avoided by choosing an appropriate frequency for the buffer mode; since the frequency gaps between storage modes are roughly equal, the buffer mode should ideally be half the gap away from the first storage mode: $\omega_{s_{1}}-\omega_{b_{1}}\approx(\omega_{s_{2}}-\omega_{s_{1}})/2$.
Wheras the swap interaction comes from terms second order in $\varphi_{c}$, the cross-Kerr/self-Kerr interaction comes from the fourth order terms. For instance, mode $B_{i}$’s self-Kerr $k_{b_{i}}$ of mode can be derived as
$U_{\text{kerr}}$ $\propto\frac{\partial^{4}U}{\partial\varphi_{c}^{4}}\Bigr|_{\varphi_{c}=0}\left(\varphi_{\text{zpf}}\frac{g_{cb_{i}}}{\Delta_{cb_{i}}}\right)^{4}(a_{cb_{i}}^{\dagger}a_{cb_{i}})^{2}$ (8)
$\propto\underbrace{(2E_{J}\cos{\varphi_{DC}})\left(\varphi_{\text{zpf}}\frac{g_{cb_{i}}}{\Delta_{cb_{i}}}\right)^{4}}_{k_{b_{i}}}(a_{cb_{i}}^{\dagger}a_{cb_{i}})^{2}$ (9)
up to a constant factor. Note that while the current discussion is limited to the coupler, the qubit also contributes to the self-kerr of the buffer modes. The cross-Kerr $\chi_{b_{i}s_{j}}$ between buffer $B_{i}$ and storage $S_{i}S_{j}$ modes can be derived in a similar way using an additional identity that the cross-Kerr is a geometric mean of the modes’ self-Kerrs:
$\chi_{b_{i}s_{j}}$ $=-2\sqrt{k_{b_{i}}k_{s_{j}}}.$ (10)
The ratio then between the undesired cross-Kerr from Eq. 10 and the desired swap interaction from Eq. 7 is
$\frac{\chi_{b_{i}s_{j}}}{g_{\text{swap},b_{i}s_{j}}}\propto\frac{\varphi_{\text{zpf}}^{2}}{\epsilon\tan{\varphi_{DC}}}\frac{g_{cb_{i}}}{\Delta_{cb_{i}}}\frac{g_{cs_{j}}}{\Delta_{cs_{j}}}$ (11)
up to a constant factor. This ratio can be optimized by reducing the zero-point phase fluctuations through arraying larger junctions in series [69]. In addition, there are other coupler designs such as LINC [41] and SNAIL [10] that completely eliminate the idle cross-Kerr between modes.
@@
The cross-Kerr between buffer mode $B_{i}$ and storage mode $S_{j}$, and the modes’ self-Kerrs can also be expressed in terms of effective mode capacitance $\{C_{b_{i}},C_{s_{j}}\}$ and inductance $\{L_{b_{i}},L_{s_{j}}\}$ as  [47]:
$k_{b_{i}(s_{j})}$ $=\frac{L_{b_{i}(s_{j})}}{C_{b_{i}(s_{j})}}\frac{C_{c}}{L_{c}}E_{C_{c}},$ (12)
@@ Appendix C Active Reset
The long coherence of the storage modes results in an extended system reset time ($5T_{1}>5$ ms). To speed up all the experiments, we implement an active reset protocol modularly, including buffer mode reset, qubit reset, and storage mode reset.
Supplementary Figure 7 (a) illustrates the sideband interactions between the buffer and dump modes. We separately prepare $\left|1\right\rangle$ in $B_{1}$ and $B_{2}$, then activate the coupler modulation at the difference frequencies of $B_{1}D_{1}$ and $B_{2}D_{2}$ for a variable time. The lossy dump modes, acting as cold reservoirs, autonomously dissipate the received excitations, effectively cooling the buffer modes. The residual population in the buffer modes is swapped back to the qubit’s $\left|f\right\rangle$ level, rotated to $\left|e\right\rangle$ level, and subsequently measured. The residual population of both buffer modes remains constant at $10\%$ after $12\,\mu\mathrm{s}$ of sideband interactions.
Supplementary Figure 7 (b) illustrates our RAQM active reset protocol, performed before the main experiments. The buffer modes are evacuated first. In the memory experiments, both the buffer modes and the qubit theoretically use, at most, the first excited states. Consequently, the protocol ensures state reset upto the second excitation for all modes. Since the qubit-$B_{i}$ dispersive shift $\chi_{qb_{i}}$ is comparable to the sideband rate, we sequentially choose $9$ different coupler RF modulations frequencies for each $B_{i}D_{i}$ sideband, accounting for the dressing effects of the dispersive shifts. After evacuating both buffer modes, we sequentially reset the qubit’s $\left|e\right\rangle$ and $\left|f\right\rangle$ states. First, we measure the qubit state using a $1\,\mu\mathrm{s}$ readout pulse, wait for $2\,\mu\mathrm{s}$, and conditionally flip the qubit state if it is in the $\left|e\right\rangle$ state. Next, we apply a qubit $\left|f\right\rangle$-$\left|e\right\rangle$ $\pi$ pulse and repeat the active reset for the $\left|e\right\rangle$ state. To reset the storage modes, we sequentially apply $7$ swap pairs, transferring the population from $S_{i}$ to $B_{1}$, then dump the population from $B_{1}$ to $D_{1}$ and autonomously evacuated. Finally, before starting the memory experiments, we idle the system for $75\,\mu\mathrm{s}$ so that any remaining low-coherence modes are fully reset. The experiments are initiated only when the qubit is confirmed in the $\left|g\right\rangle$. The qubit readout performance in our experiments is shown in Supplementary Table 4.
[…]
@@ Appendix D System Temperature
We benchmark the cavity’s temperature using the Ramsey parity measurement [16]. To detect the thermal photon population in $B_{1}$, we repeatedly measure the parity information 80 times in a single experiment. Each measurement sequence consists of the following pulses: a transmon $\pi/2$ rotation, an idle period of $\pi/\chi$, another transmon $\pi/2$ rotation, a readout pulse ($1\,\mu\mathrm{s}$), and a readout evacuation period ($2.5\,\mu\mathrm{s}$). Each experiment produces an 80-bit string indicating the transmon’s state. Supplementary Figure 8(a) illustrates the string behavior when $B_{1}$ is initialized to $\left|1\right\rangle$.
[…]
@@
In Supplementary Figure 8(b), we analyzed $10^{5}$ measurements taken when the system is in its thermal equilibrium state. A higher ratio in the plot indicates a greater likelihood of photon population in the buffer modes. As the threshold increases (dashed line), measurement errors are filtered out, isolating events corresponding to thermal photons.
After extracting the thermal photon events, we need to compensate for two key factors to avoid overestimating the system temperature: detection efficiency and event assignment. The detection efficiency $\eta_{i}$ describes the probability of detecting a photon when $B_{i}$ is in $\left|1\right\rangle$. In our experiments, we separately prepare $\left|1\right\rangle$ in $B_{1}$ and $B_{2}$, perform the Ramsey parity measurement to count cavity photon events, and approximate $\eta_{i}$ using $2\times 10^{5}$ experiments. In Supplementary Figure 8(c), fewer photon events are captured as the threshold increases. Since both $B_{1}$ and $B_{2}$ are coupled to the transmon, thermal population in either cavity can trigger the detection. We assume that both modes have the same temperature $T_{c}$. Given the experimentally measured total photon counts $\delta$ and efficiency $\eta_{i}$, the temperature $T_{c}$ required to explain the counts should satisfy:
$\sum_{i=1,2}\frac{\eta_{i}}{-1+\exp{(\frac{\hbar\omega_{bi}}{k_{B}T_{c}}})}=\delta\$(14)
[…]
@@ Appendix E Phase Calibration for RAQM experiments
Figure 9: RAQM Phase calibration. Each transmon gate’s virtual-$Z$ phase comprises three components: the active-access phase $\theta_{ii}$, the inactive-access phase $\theta_{ij}$ accumulated during storage read and write, and the time-dependent idle phase $\theta_{\Delta_{i}}(t)$ caused by hardware frequency tracking. (a) Pulse sequences for state transfer between transmon and $B_{1}$. (b) active-access phase $\theta_{ii}$ calibration for $S_{i}$. (c) Idle phase calibration extracts the Stark-shift $\Delta_{i}$ to $S_{i}$ through the Ramsey-like experiments. (d) Inactive-access phase $\theta_{ij}$ calibration on $S_{i}$ after $S_{j}$ access operations. $\theta_{ij}$ are calculated based on the maximum contrast phase $\tilde{\theta}_{i}j$ and $\Delta_{i}$. To improve calibration accuracy, $N+1$ writes and NN reads are repeated.
@@
Supplementary Figure 9 shows the process of virtual-$Z$ phase calibration. For simplicity, state read and write operations between the transmon and $B_{1}$ are represented using circuit symbols in Supplementary Figure 9(a). Supplementary Figure 9(b) shows the circuit to calibrate $\theta_{ii}$. The phase of the second transmon gate is swept, and $\theta_{ii}$ is determined as the phase that maximizes the final $\left|e\right\rangle$ population. Supplementary Figure 9(c) shows the circuit to calibrate $\theta_{ii}$. By sweeping the waiting time between two $B_{1}S_{i}$ swaps, we fit the Ramsey signal and extract the Stark-shift frequency $\Delta_{i}$. Supplementary Figure 9(d) shows the circuit to calibrate $\theta_{ij}$. Since the inactive-access phase is small, we repeat the read and write operations on $S_{j}$ for $5$ times. The maximum contrast phase $\tilde{\theta}_{ij}$, determined by the maximum final $\left|e\right\rangle$ population in the transmon, can be used to calculate $\theta_{ij}$: